If n is any even integer and m is any odd integer, then (n + 2)2 - (m - 1)2 is even.
1
Expert's answer
2020-10-19T18:11:24-0400

Since "n" is even integer and "m" is odd integer, then "n=2k" and "m=2s+1" for some "k,s\in \mathbb Z".

Then "(n+2)^2-(m-1)^2= (2k+2)^2-(2s+1-1)^2=2^2(k+1)^2-(2s)^2=4(k+1)^2-4s^2=4((k+1)^2-s^2)" where "(k+1)^2-s^2" is integer. Therefore, "(n+2)^2-(m-1)^2" is even.